OpenStudy (kj4uts):

Reflecting the graph y=sin x across the y-axis is the same as reflecting it across the x-axis. True or False. I think it would be true because it you rotated the graph it would still match up. Please explain. Thank you!

OpenStudy (anonymous):

What do you do to a function to reflect across the x-axis? and the y-axis?

OpenStudy (aum):

y = f(x) f(-x) reflects it across y-axis. -f(x) reflects it across x-axis.

OpenStudy (anonymous):

So mathematically you'd need to do different things to the graph to reflect it either the x or the y, I think the results are the same though.

OpenStudy (aum):

y = sin(x) --- (1) y = sin(-x) reflects it across y-axis. --- (2) y = -sin(x) reflects it across x-axis. --- (3) Since sin(-x) = -sin(x), operations (2) and (3) have the same effect.